OpenStudy (anonymous):

Solve for R. C=1/9π n^2r

OpenStudy (abb0t):

For this, you are going to use natural log to solve for r. I am going to assume that your function looks like this: \[\frac{ 1 }{ 9 }\pi n^{2r} = C\] Follow these steps: 1. multiply 2. divide 3. take the natural log [NOTE: \[\ln(a)^P=pln(a)\] From this point, you should be able to solve use the algebra tricks similar to the previous steps to solve for 'r'. Let me know if you get stuck.

OpenStudy (anonymous):

The r is suppose to be regular not up by the small 2

OpenStudy (anonymous):

is that the question?\[C=\frac{1}{9}π n^2r\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[C=\frac{1}{9}π n^2r\] multiply both by 9 \[9*C=\frac{1}{9}π n^2r*9\] \[9*C=π n^2r\] divide both sides by π n^2 \[\frac{1}{π n^2}9C=\frac{1}{π n^2}π n^2r\] \[\frac{1}{π n^2}9C=r\] \[\frac{9C}{π n^2}=r\]